Circular No. 6354

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) J. Lecacheux, Meudon Observatory; L. Jorda, Meudon Observatory

and Laboratoire d’Astronomie Spatiale, Marseille; A. Enzian, Laboratoire de Glaciologie, Université de Grenoble; J. Klinger, Laboratoire de Glaciologie, Grenoble; F. Colas, Bureau des Longitudes, Paris; E. Frappa, Planetarium of Saint-Etienne; and P. Laques, Pic du Midi Observatory, write: “Circumnuclear dust jets of the comet have become an easy target for telescopes equipped with accurate guiding systems. Filters selecting the dust continuum (such as those covering 484.5–486.5 nm and 684–689 nm in the IAU standard set) provide better contrast, but broadband filters permit shorter exposures. The synodic period of rotation of the nucleus prior to Mar. 23 was found to be a submultiple of 24 hr, as main jets crossed the nucleus-sun line each day at the same hour. At present our preferred period is 6.0 hr (revising our estimate on IAUC 6344), but we do not exclude the value of 8.0 hr. Cooperative coverage by observatories at various longitudes is needed. For several days following Mar. 23, the viewing geometry of the nucleus for earth-based observers was rapidly changing. The inner part of the tail of C/1996 B2 (∼ 5000 km long) has a conspicuous yellow color to telescopic visual inspection. From our multicolor CCD imaging with the Pic-du-Midi 1.05-m reflector, it clearly has the photometric properties of dust reflecting sunlight. We have used, since Mar. 10, a set of 11 narrow-band filters (in the range 426–925 nm), including CO^{+} and H_{2}O^{+} emissions, C_{2} emission and adjacent continuum, plus various broadband red filters for 830 nm and longer wavelengths. The dust particles have been collimated in a narrow region of apparent width 300 km in the exact antisolar direction during Mar. 10–23. During Mar. 23.9–24.2 UT, we detected luminous knots slowly receding from the nucleus. The faintest, of mag 13, was at the projected distance of 21′′.5 on Mar. 24.0 and moved outwards at the projected velocity of 0′′.8/hr; assuming that this object followed the extended solar radius vector, the geometric conditions at this time imply a real distance from the nucleus of 2340 km and an average real recession velocity of 12 m/s. A second, more luminous knot was closer to the nucleus and receded similarly. The luminosity of the inner tail, especially its 2000-km inner part, increased conspicuously between Mar. 21.2 and 21.9.”

1996 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green